OpenStudy (anonymous):

which points have a slope of 1/3? (3, 2) (9, 4) (2, 3) (9, 4) (2, 3) (4, 9) (3, 2) (4, 9)

jimthompson5910 (jim_thompson5910):

Hint: use the slope formula m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

do you know it though?

jimthompson5910 (jim_thompson5910):

so for instance, choice D has a slope of... m = (y2-y1)/(x2-x1) m = (9-2)/(4-3) m = 7/1 m = 7

jimthompson5910 (jim_thompson5910):

So choice D does NOT have a slope of 1/3

OpenStudy (anonymous):

okay then.

jimthompson5910 (jim_thompson5910):

see how to get the answer?

OpenStudy (anonymous):

yes, thank you

jimthompson5910 (jim_thompson5910):

you're welcome